2016 · Q85
There are four identically sized and shaped balls in a box, with only its top open. Each ball is of a different colour, these being: Green, Red, White and Blue, only one ball of each colour. Without looking into the box, one ball is randomly picked out and its colour is noted; then it is returned to the box. What are the chances that, in two successive draws, one may get the white ball and the red ball – in whichever order?
General Mental Ability, Quantitative Aptitude and StatisticsElementary statistics
- A1/16
- B1/8
- C1/4
- D1/2
Explanation
With replacement, P(white then red) = 1/4 × 1/4 = 1/16 and P(red then white) = 1/16. Either order gives 1/16 + 1/16 = 2/16 = 1/8. Hence (B).